Solve d^{2}y/dx^{2}+2dy/dx+y=xsinx


1
Expert's answer
2021-08-24T06:18:07-0400
"y''(x)+2y'(x)+y(x)=x\sin x"

The characteristic equation of the homogeneous DE

"k^2+2k+1=0"

Roots

"k_1=k_2=-1"

Hence, the general solution of homogeneous DE

"y_h(x)=C_1e^{-x}+C_2xe^{-x}"

The partial solution of inhomogeneous DE

"y_i(x)=(A+Bx)\sin x+(C+Dx)\cos x"

We get

"y'_i(x)=(A+Bx)\cos x-(C+Dx)\sin x\\+B\sin x+D\cos x"

"y''_i(x)=-(A+Bx)\sin x-(C+Dx)\cos x\\+2B\cos x-2D\sin x"

"-(A+Bx)\sin x-(C+Dx)\cos x\\+2B\cos x-2D\sin x+2((A+Bx)\cos x\\-(C+Dx)\sin x+B\sin x+D\cos x)\\+(A+Bx)\sin x+(C+Dx)\cos x\\=x \sin x"

So, the partial solution of inhomogeneous DE

"y_i(x)=\frac{1}{2}(\cos x-x\cos x+\sin x)"

Finally,

"y(x)=C_1e^{-x}+C_2xe^{-x}\\+\frac{1}{2}(\cos x-x\cos x+\sin x)"